A father is four times as old as his daughter. In 7 years, he will be three times as old as she is. How old is the daughter now?
Answer by richwmiller(17219)

f=4d
f+7=3*(d+7)
4d+7=3d+21
d=14
f=4*14=56
check
56+7=3*(14+7)
63=3*21
63=63
ok